OpenStudy (anonymous):

\[(x^6 y ^{-3} )^\frac{ 1 }{ 2 } y^2/\sqrt[3]{x}=x^p y^q\] HELP!

OpenStudy (popitree):

\[={ (x ^{3}y ^{-1})y ^{2}}/x ^{1/3}\]

OpenStudy (popitree):

\[= (x ^{3}y)x ^{-1/3}\]

OpenStudy (popitree):

=y/x

OpenStudy (popitree):

p=-1, q=1

OpenStudy (anonymous):

i got lost in the second step, how did u get rid of the division and the -1 exponent?

OpenStudy (popitree):

y^{-1}*y^{2}

OpenStudy (popitree):

this one?

OpenStudy (anonymous):

oh i see it, that was it!

OpenStudy (popitree):

so u got the ans.. any confusion

OpenStudy (anonymous):

arent u suppose to multiply y^2 with X^3 since x^3 is also inside the brackets?

OpenStudy (popitree):

that will happen during addition or subtraction

OpenStudy (popitree):

(4*5)*2 = 8*5 = 4*20 :)

OpenStudy (popitree):

4*10 .. typing mistake

OpenStudy (popitree):

it is not 8*10

OpenStudy (anonymous):

oh ok, thnx!

OpenStudy (popitree):

cool, no problem